Suppose we wish to mix some 32% salt solution with some 58% salt solution to produce ten liters of 46.3% salt solution. How many liters of each of the starting solutions needs to be used?
Answer by checkley77(9413)

.58x+.32(10-x)=.463*10
.58x+3.2-.32x=4.63
.26x=4.63-3.2
.26x=1.43
x=1.43/.26
x=5.5 liters of 58% solution is used.
10-5.5=4.5 liters of 32% solution is used.
Proof:
.58*5.5+.32*4.5=.463*10
3.19+1.44=4.63
4.63=4.63